Island Novelties, Inc., of Palau makes two products—Hawaiian Fantasy and Tahitian Joy. Each product’s selling price, variable expense per unit, and annual sales volume are as follows:

Hawaiian Fantasy

Tahitian Joy

Selling price per unit

$

20

$

125

Variable expense per unit

$

13

$

50

Number of units sold annually

15,000

5,100

Fixed expenses total $325,000 per year.

Required:

1. Assuming the sales mix given above, do the following:

a. Prepare a contribution format income statement showing both dollar and percent columns for each product and for the company as a whole.

b. Compute the company's break-even point in dollar sales. Also, compute its margin of safety in dollars and its margin of safety percentage.

2. The company has developed a new product called Samoan Delight that sells for $45 each and that has variable expenses of $27 per unit. If the company can sell 12,500 units of Samoan Delight without incurring any additional fixed expenses:

a. Prepare a revised contribution format income statement that includes Samoan Delight. Assume that sales of the other two products does not change.

b. Compute the company’s revised break-even point in dollar sales. Also, compute its revised margin of safety in dollars and margin of safety percentage.
Answer #1
1-a.
Island Novelties, Inc.
Contribution Income Statement
Hawaiian Fantasy Tahitian Joy Total
Amount % Amount % Amount %
Sales 300000 100% 637500 100% 937500 100%
Variable expenses 195000 65% 255000 40% 450000 48%
Contribution margin 105000 35% 382500 60% 487500 52%
Fixed expenses 325000
Net operating income 162500
1-b.
Break-even point in dollars = Fixed expenses / Contribution margin ratio of company = 325000 / 52% 625000
Margin of safety in dollars = Sales - Break even point in dollars = 937500 - 625000 312500
Margin of safety percentage = Margin of safety in dollars / Sales = 312500 / 937500 33.3%
2-a.
Island Novelties, Inc.
Contribution Income Statement
Hawaiian Fantasy Tahitian Joy Samoan Delight Total
Amount % Amount % Amount % Amount %
Sales 300000 100.0% 637500 100.0% 562500 100.0% 1500000 100.0%
Variable expenses 195000 65.0% 255000 40.0% 337500 60.0% 787500 52.5%
Contribution margin 105000 35.0% 382500 60.0% 225000 40.0% 712500 47.5%
Fixed expenses 325000
Net operating income 387500
2-b.
Break-even point in dollars = Fixed expenses / Contribution margin ratio of company = 325000 / 47.5% 684211
Margin of safety in dollars = Sales - Break even point in dollars = 1500000 - 684211 815789
Margin of safety percentage = Margin of safety in dollars / Sales = 815789 / 1500000 54.4%